On January 1, 2017, Bobcat Company sold 12% bonds having a face value of $300,000 to yield 10% market rate. The bonds are dated January 1, 2017 and mature January 1, 2022 (5 year term), with interest payable quarterly (April 1, July 1, October 1, and January 1) each year. Instructions Prepare complete bond amortization schedules for Bobcat using the Excel templates on the following tabs. Tab 1: Assume Bobcat allocates interest and unamortized discount or premium on the EFFECTIVE-INTEREST (EI) basis. Tab 2: Assume Bobcat allocates interest and unamortized discount or premium on the STRAIGHT-LINE (SL) basis.

Effective Interest Basis

Date Payment Number

Cash Interest Expense

Interest Expense Premium Amortization Amount

Carrying Amount of Bonds

1/1/2017 Start
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20

Straight-Line

Date

Payment

Number

Cash

Interest

Expense

Interest

Expense

Premium

Amortization

Amount

Carrying

Amount of

Bonds

start
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20

Please explain excel formulas used or how you solved.
Answer #1
Ans 1.
Bobcat Company
Bond Face value $                300,000
Coupon Rate = 12%
Quarterly Interest payment = $                    9,000
Market Yield =Effective Interest 10%
Quarterly Market yield =Effective Interest= 2.50%
Maturity period in years                                5
Total Quarterly period for coupons payment                              20
PV Annuity Factor for 20 Qtrs @2.5% =[1-(1.025)^-20]/2.5%=                 15.58916
PV factor after 20 Qtrs @2.5% =1/1.025^20=                    0.61027
Current Bond Price calculation
Cash Flow Amt PVIF/PV Factor PV of Cash flows
Quarterly Interest payments                        9,000                  15.58916 $ 140,302.44
Marturity Value                    300,000                    0.61027 $ 183,081.00
Total PV of Cash flows=Bond Current Price $ 323,383.44
So Bond Sales Price $          323,383.44
Bond Par Value = $          300,000.00
Premium $            23,383.44
Premium Amortization by Effective Interest rate a b c d=b-c e= prev prem bal-d =300,000+e
Payment period Opening Bond Payable Amount Cash Interest Exp Interest Exp @2.5% Premium Amortization Balance Premium Carrying Bond Amt
start: 01/01/2017                    323,383 $                     9,000           8,084.59 $           915.41 $     22,468.03 $ 322,468.03
2                    322,468 $                     9,000           8,061.70 $           938.30 $     21,529.73 $ 321,529.73
3                    321,530 $                     9,000           8,038.24 $           961.76 $     20,567.97 $ 320,567.97
4                    320,568 $                     9,000           8,014.20 $           985.80 $     19,582.17 $ 319,582.17
5                    319,582 $                     9,000           7,989.55 $       1,010.45 $     18,571.72 $ 318,571.72
6                    318,572 $                     9,000           7,964.29 $       1,035.71 $     17,536.02 $ 317,536.02
7                    317,536 $                     9,000           7,938.40 $       1,061.60 $     16,474.42 $ 316,474.42
8                    316,474 $                     9,000           7,911.86 $       1,088.14 $     15,386.28 $ 315,386.28
9                    315,386 $                     9,000           7,884.66 $       1,115.34 $     14,270.93 $ 314,270.93
10                    314,271 $                     9,000           7,856.77 $       1,143.23 $     13,127.71 $ 313,127.71
11                    313,128 $                     9,000           7,828.19 $       1,171.81 $     11,955.90 $ 311,955.90
12                    311,956 $                     9,000           7,798.90 $       1,201.10 $     10,754.80 $ 310,754.80
13                    310,755 $                     9,000           7,768.87 $       1,231.13 $       9,523.67 $ 309,523.67
14                    309,524 $                     9,000           7,738.09 $       1,261.91 $       8,261.76 $ 308,261.76
15                    308,262 $                     9,000           7,706.54 $       1,293.46 $       6,968.30 $ 306,968.30
16                    306,968 $                     9,000           7,674.21 $       1,325.79 $       5,642.51 $ 305,642.51
17                    305,643 $                     9,000           7,641.06 $       1,358.94 $       4,283.57 $ 304,283.57
18                    304,284 $                     9,000           7,607.09 $       1,392.91 $       2,890.66 $ 302,890.66
19                    302,891 $                     9,000           7,572.27 $       1,427.73 $       1,462.93 $ 301,462.93
20                    301,463 $                     9,000           7,536.57 $       1,463.43 $ 300,000.00
Ans 2.
So Bond Sales Price $          323,383.44
Bond Par Value = $          300,000.00
Premium $            23,383.44
Quarterly Premium Amortization over 20 Qtrs $              1,169.17
Premium Amortization by SL Amortization a b c=fixed amt d=b-c e= prev bal prem-c =$300,000+e
Payment period Opening Bond Payable Amount Cash Interest Exp Premium Amortization Interst Expense Balance Premium Carrying Bond Amt
start: 01/01/2017                    300,000 $                     9,000 $      1,169.17 $       7,830.83 $     22,214.27 $ 322,214.27
2                    322,214 $                     9,000 $      1,169.17 $       7,830.83 $     21,045.10